OpenStudy (gabrielah96):

two small frogs leap straight up from a lily pad. frog A leaps with an initial velocity of 0.451m/s while frog B leaps with an initial velocity of 1.88 m/s. when from A lands back on the lily pad what is the position and velocity of frog B? take upwards to be positive and let the position of the lily pad be zero

OpenStudy (phi):

do you know that V= a * t here , acceleration is due to gravity -9.8 m/s^2 (i.e. downward) I would find the time it takes for the velocity to slow to 0 (that will be at its peak before falling back down). Because this will be same time as falling, we could do: -0.451 m/s = -9.8 m/s^2 t t = 0.451/9.8 now double that to get the total time up and back down. what do you get ?